OpenStudy (anonymous):

What is the simplified form of sin(x + p)? a. cos x b. sin x c. –sin x d. –cos x

OpenStudy (anonymous):

again go with option C because \[\sin(x+\pi)=-\sin(x)\]

OpenStudy (anonymous):

in the question what does P stand for?

OpenStudy (anonymous):



OpenStudy (anonymous):

pretty sure from the question that it mean \(\pi\)

OpenStudy (anonymous):

ohh , okay i understand that part now